Eight boys and three girls are to sit in a row for a photograph. Find the probability that no two girls are together.


1
Expert's answer
2021-01-08T15:27:22-0500

Given: "8" boys; "3" girls;

total people: "11"

total ways of arranging: "11!"

We can place the boys in 8! "= P^8_8" ways and the girls can't sit next to each other,

there are only 9 seats for the girls, but there are only 3 girls hence there are only "P^9_3" ways for them to sit.


All posible ways of arranging which are suitable for us: "P^8_8*P^9_3 = \frac{8!}{0!}*\frac{9!}{(9-3)!}, 0! =1"

"p =\frac{P^8_8*P^9_3}{11!} = \frac{8!}{0!}*\frac{9!}{(9-3)!11!} \approx 0.509091"

"p \approx 0.509091"